# On distinguishing coloring and some variants of proper coloring of graphs derived from subdivision operations

### v1

@@ Abstract.
Let $G$ be a simple, finite, connected, and undirected graph, and $T$ be a finite tree. The middle graph $M(G)$ of $G$ is obtained from the subdivision graph $S(G)$ after joining pairs of subdivided vertices that lie on adjacent edges of $G$ and the central graph $C(G)$ of $G$ is obtained from $S(G)$ after joining all non-adjacent vertices of $G$. We show that if the order of $G$ is at least $4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$(as abstract groups) and apply these results to obtain new sharp upper bounds of the distinguishing number and the distinguishing index of $C(G)$ and $M(G)$ inspired by an algorithm due to Kalinowski, Pilśniak, and Woźniak from 2016. Furthermore, we study the total distinguishing chromatic number of $C(G)$ and $S(G)$, use Latin squares to verify the AVD-total coloring conjecture for central graphs of regular graphs and some other classes of graphs (which is a partial progress towards answering an open question of Panda, Verma, and Keerti), and obtain new bounds of the total dominator chromatic number of $C(G)$ and $C(T)$.
@@
Automorphism groups, distinguishing coloring, total dominator coloring, adjacent vertex distinguishing total coloring, central graph, subdivision graph
@@
Corresponding author.
@@ 1. Introduction
Let $Aut(G)$ be the automorphism group of a graph $G=(V(G),E(G))$, $id_{G}$ be the trivial automorphism of $G$, and $f$ be a vertex coloring or an edge coloring or a total coloring (that is, a coloring of its vertices and edges) of $G$. An automorphism $\phi$ of $G$ preserves $f$ if the following holds: (1) $f(x)=f(\phi(x))$ for all $x\in V(G)$ if $f$ is a vertex coloring, (2) $f(e)=f(\phi(e))$ for all $e\in E(G)$ if $f$ is an edge coloring, (3) $f(x)=f(\phi(x))$ for all $x\in V(G)\cup E(G)$ if $f$ is a total coloring.
1.1. Distinguishing number and distinguishing index
Albertson and Collins [1] introduced the distinguishing number $D(G)$ of a graph $G$ as the least integer $d$ such that $G$ admits a vertex coloring with $d$ colors that is only preserved by $id_{G}$ and Kalinowski and Pilśniak [16] introduced the distinguishing index $D^{\prime}(G)$ of $G$ as the least integer $d$ such that $G$ admits an edge coloring with $d$ colors that is only preserved by $id_{G}$. In [4, Theorems 3.8 and 4.3], Alikhani and Soltani proved that if $G$ is a connected graph of order $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$ where $\Delta(G)$ is the maximum degree of $G$. Hamada and Yoshimura [13] and Vernold [31] introduced the concept of middle graphs and central graphs, respectively; classes of graphs derived from subdivision graphs. Since then, the notion of central graphs and middle graphs has been intensively studied in graph theory; see, for example, [5, 9, 11, 12, 18, 19, 23, 28].
Results: Let $G$ be a connected graph of order $n$. The following holds:
@@ 1.1. Distinguishing number and distinguishing index
    $D^{\prime}(M(G))\leq 3$ for any positive integer $n$, and
@@
    If $n\geq 4$, then $D^{\prime}(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$, and $D(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
Moreover, the bounds mentioned in (1) and (3) are sharp (see Theorems 3.2, 3.4, 3.6).
1.2. Total distinguishing chromatic number
A proper total coloring of a graph $G$ is a total coloring of $G$ so that no adjacent vertices, no adjacent edges, and no incident vertices and edges get the same color. The least number of colors needed for a proper total coloring of $G$ is called its total chromatic number, denoted by $\chi^{\prime\prime}(G)$.
Behzad [6] and Vizing [33] independently posed the following conjecture.
Conjecture 1.1.
(Total-Coloring Conjecture (TCC)). For any graph $G$, $\chi^{\prime\prime}(G)\leq\Delta(G)+2$.
Recently, Panda et al. [28, Theorem 3.1] proved that TCC holds for the central graph of any graph of order $n\geq 5$. We observe an alternative short proof of the result using the following lemma.
Lemma 1.2.
([36, Lemma 2.4]) If $G$ is a graph of order $n$ that contains an independent set $S$ such that $|S|\leq n-\Delta(G)-1$, then $G$ satisfies TCC.
Theorem 1.3.
(Panda–Verma–Keerti;[28]) TCC holds for the central graph of any graph.
Short proof of Theorem 1.3.
Let $C(G)$ be the central graph of a graph $G=(V(G),E(G))$ where $V_{1}=V(G)$, $V_{2}=V(C(G))\backslash V(G)$, $|V_{1}|=n$, and $|V_{2}|=r$. If $n\geq 3$, then the order of $C(G)$ is $n+r$, $\Delta(C(G))=n-1$, and $C(G)$ contains an independent set $V_{2}$ of size $r=(n+r)-\Delta(C(G))-1$. By Lemma 1.2, $C(G)$ satisfies TCC. If $n\leq 2$, then it is trivial to verify. ∎
We say $G$ is of Type $j$ if $\chi^{\prime\prime}(G)=\Delta(G)+j$ for $j\in\{1,2\}$. In [15], Kalinowski, Pilśniak, and Woźniak introduced the total distinguishing chromatic number $\chi^{\prime\prime}_{D}(G)$ of a graph $G$ as the least integer $d$ such that $G$ has a proper total coloring with $d$ colors that is preserved only by $id_{G}$. We study the total distinguishing chromatic number of central and subdivision graphs.
Results: Let $G$ be a connected graph of order $n$. The following holds:
@@ 1.2. Total distinguishing chromatic number
    (Proposition 4.4) $\Delta(C(G))+1\leq\chi^{\prime\prime}(C(G))\leq\chi^{\prime\prime}_{D}(C(G))\leq\Delta(C(G))+2$.
@@
    (Theorem 4.5(1)) If $G$ is regular and $n\geq 5$ is odd, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
@@
    (Theorem 4.5(2)) If $G$ is regular and a non complete graph, $n\geq 5$ is even, and $\Delta(G)\leq\frac{2n-29}{6}$, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
@@
    (Proposition 4.9) $\chi^{\prime\prime}_{D}(S(G))$ is either $\Delta(S(G))+1$ or $\Delta(S(G))+2$.
1.3. AVD-total chromatic number
An adjacent vertex distinguishing (AVD) total coloring of a graph $G$ with edge set $E(G)$ is a proper total coloring $f$ such that for any edge $\{u,v\}\in E(G)$, we have $C_{G}(u)\neq C_{G}(v)$, where $C_{G}(u)=\{f(u)\}\cup\{f(\{u,v\}):\{u,v\}\in E(G)\}$ is the color class of $u$. The adjacent vertex distinguishing (AVD)-total chromatic number $\chi^{\prime\prime}_{a}(G)$ of $G$ is the minimum integer $k$ such that there exists an AVD-total coloring of $G$ using $k$ colors. Zhang et al. [37] introduced AVD-total-colorings and posed the following conjecture.
Conjecture 1.4.
(AVD-TCC). For any graph $G$, $\chi^{\prime\prime}_{a}(G)\leq\Delta(G)+3$.
We say $G$ is of AVD-Type $i$ if $\chi^{\prime\prime}_{a}(G)=\Delta(G)+i$ for $i\in\{1,2,3\}$. The Conjecture 1.4 for general graphs $G$ with $\Delta(G)=3$ has been verified by Wang [34] and Chen [7]. Lu et al. [22] validated the conjecture for all graphs $G$ with $\Delta(G)=4$. The Conjecture 1.4 is also known to be true for many families of graphs such as 4-regular graphs [26], hypercubes [8], complete equipartite graphs [24], indifference graphs [27], split graphs [32], and several other classes of graphs. Panda et al. [28] obtained the AVD-total chromatic number of central graphs for some known families of graphs and mentioned the following open question (see [28, Section 5]).
Question 1.5.
(Panda et al. [28]) Does AVD-TCC hold for the central graph of any graph?
In [28, Section 5], it was mentioned that it would be interesting to solve the AVD-total coloring classification problem on the central graph of any graph. We have made partial progress in this direction.
Results: The following holds:
@@ 1.3. AVD-total chromatic number
    (Theorem 5.1) If $G$ is a regular graph of even order $n\geq 5$, then $C(G)$ is AVD-Type 2 and if $G$ is a regular graph of order $n\geq 5$, then $C(G)$ satisfies AVD-TCC.
@@
    (Theorem 5.3) If $\Delta(G)\geq 5$, then the subdivision graph $S(G)$ of $G$ is AVD-Type 1.
  3. (3)
    (Proposition 5.5 (4)) If two graphs $G_{1}$ and $G_{2}$ have the same order and $G_{1}+G_{2}$ is the join of $G_{1}$ and $G_{2}$, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
1.4. Total dominator chromatic number
The notion of total domination number was introduced by Cockayne, Dawes and Hedetniemi in [10] and has been studied for central graphs and middle graphs, see [9, 18, 19, 20, 21]. In particular, Kazemnejad and Moradi [19] and Chen et al. [9] studied sharp upper bounds for the total domination number of central graphs of connected graphs and trees. Motivated by the relation between the chromatic number and the total domination number, the notion of the total dominator chromatic number was studied in [17]. A total dominator coloring, briefly TDC, of a graph $G$ is a proper vertex coloring of $G$ in which each vertex of $G$ is adjacent to every vertex of some color class. The total dominator chromatic number $\chi^{t}_{d}(G)$ of $G$ is the minimum number of color classes in a TDC of $G$. The total dominator chromatic number was studied for middle graphs by Kazemnejad et al. [20], $k$-subdivision of graphs by Alikhani et al. [2], and central graphs by Kazemnejad et al. [18]. In [18, Theorem 2.1], Kazemnejad and Kazemi proved that the following bounds are sharp if $G$ is a connected graph of order $n\geq 2$:
  • $\lfloor 2n/3\rfloor+1\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+2$ if $\Delta(G)\leq n-2$.
  • $\chi^{t}_{d}(C(G))\leq n+\lceil n/2\rceil$ if $G$ has a Hamiltonian path.
Results: Let $G$ be a connected graph of order $n\geq 5$. The following holds:
  1. (1)
    If $\Delta(G)\leq n-3$ and $\overline{G}$ is the complement of $G$, then $\chi^{t}_{d}(\overline{G})\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
  2. (2)
    If $G$ is a tree, then $\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
Moreover, the upper bounds of $\chi^{t}_{d}(C(G))$ mentioned in (1) and (2) are sharp (see Theorem 6.1).
1.5. Tables of results
Let $C_{n}$ be a cycle graph of $n$ vertices, $P_{n}$ be a path of $n$ vertices, $K_{n}$ be a complete graph of $n$ vertices, and $K_{n,m}$ be a complete bipartite graph with bipartitions of sizes $n$ and $m$. A claw is the graph $K_{1,3}$. The degree of a vertex $v\in V(G)$ in $G=(V(G),E(G))$, denoted by $deg_{G}(v)$, is the number of edges that emerge from $v$. We say $G$ is a regular graph if all vertices of $G$ have the same degree. The degree of a regular graph $G$, denoted by $d(G)$, is the same as the maximum degree $\Delta(G)$ and the minimum degree $\delta(G)$. Let $N_{G}(v)=\{u\in V(G):\{u,v\}\in E(G)\}$ for any $v\in V(G)$.
Graph $D^{\prime}(H)$ $D(H)$
$H=C(G),n\geq 4$ $\leq\lceil\sqrt{\Delta(G)}\rceil$ (Thm. 3.2) $\leq\lceil\sqrt{\Delta(G)}\rceil$ (Thm. 3.4)
$H=M(G),n\geq 3$ $\leq 3$ (Thm. 3.6(2)) $\leq\Delta(G)$ (Thm. 3.6(1))
$H=S(G)$ $\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$ ([4, Thm. 4.3]) $\leq\lceil\sqrt{\Delta(G)}\rceil$ ([4, Thm. 3.8])
Graph $\chi^{\prime\prime}_{a}(H)$ $\chi^{\prime\prime}_{D}(H)$
$H=S(G),\Delta(G)\geq 5$ $\Delta(H)+1$ (Thm. 5.3) $\leq\Delta(H)+2$ (Prop. 4.9)
$H=C(G),n\equiv 0$(mod 2), $G\not\cong K_{n}\text{ is regular}$, $n\geq 5$, and $\Delta(G)\leq\frac{2n-29}{6}$ $\Delta(H)+2$ (Thm. 5.1) $\Delta(H)+1$ (Thm. 4.5)
[…]
$H=C(K_{m,r})$ $\leq\Delta(H)+3$ (Prop. 5.5(2)) $\leq\Delta(H)+2$ (Prop. 4.4)
$H=C(G_{1}+G_{2})$, $G_{1},G_{2}\in\{P_{n},C_{n},K_{n},K_{n,r}\}$ $\leq\Delta(H)+3$ (Prop. 5.5(3)) $\leq\Delta(H)+2$ (Prop. 4.4)
$H=C(H_{1}+H_{2})$, order of $H_{1}$ = order of $H_{2}$ $\leq\Delta(H)+3$ (Prop. 5.5(4)) $\leq\Delta(H)+2$ (Prop. 4.4)
Graph $\chi^{t}_{d}(H)$
$H=C(G),n\geq 5,\Delta(G)\leq n-3$ $\geq\chi^{t}_{d}(\overline{G})$, $\leq\Delta(H)+1$ (Thm. 6.1(1))
$H=C(T),T\text{ is a tree, }\text{order of }T\geq 5$ $\leq\Delta(H)+1$ (Thm. 6.1(2))
$H=M(G),n\geq 2,G\text{ has $m$ edges}$ $\leq n+m-1$ ([20, Thm. 2.3])
$H=S(G),G\text{ has $m$ edges}$ $\leq 2m$ ([2, Thm. 2.2])
$H=C(G),n\geq 3,\Delta(G)\leq n-2$ $\leq\Delta(H)+2$ ([18, Thm. 2.1])
$H=C(G),n\geq 2,\text{ $G$ has a Hamiltonian path}$ $\leq n+\lceil n/2\rceil$ ([18, Thm. 2.1])
In the above tables, we list the results obtained in this manuscript, where $n$ denotes the order of $G$.
@@ 2. Automorphism groups
Definition 2.1.
The subdivision graph $S(G)$ of a graph $G=(V(G),E(G))$ is a bipartite graph with the set of vertices $V(S(G))=V_{1}\cup V_{2}$ where $V_{2}=\{w_{v,v^{\prime}}:w_{v,v^{\prime}}\text{ subdivide the edge }\{v,v^{\prime}\}\in E(G)\}$ and $V_{1}=V(G)$, and the set of edges $E(S(G))=\{\{v,w_{v,v^{\prime}}\},\{w_{v,v^{\prime}},v^{\prime}\}\}:\{v,v^{\prime}\}\in E(G)\}$.
Definition 2.2.
The central graph $C(G)$ of a graph $G=(V(G),E(G))$ of order $n$ and size $m$ is a graph of order $n+m$ and size ($n\choose{2}$ + $m$ that is obtained by subdividing each edge of $G$ exactly once and joining all the non-adjacent vertices of $G$ in $C(G)$. In particular, $C(G)$ is a graph with the set of vertices $V(C(G))=V(S(G))$ and the set of edges $E(C(G))=E(\overline{G})\cup\{\{v,w_{v,v^{\prime}}\},\{w_{v,v^{\prime}},v^{\prime}\}\}:\{v,v^{\prime}\}\in E(G)\}$ where $\overline{G}$ is the complement of $G$.
In this section, we sketch some lemmas concerning $Aut(C(G))$ and $Aut(M(G))$ where $M(G)$ is the middle graph of $G$. Applying these lemmas, we will prove Theorems 3.2, 3.4, and 4.5.
@@
First, we note that $N_{C(G)}(a)\neq N_{C(G)}(b)$ for any $a,b\in V_{2}$. Let $x\in V_{2}$. If $\psi(x)=y$ for some $y\in V_{2}\backslash\{x\}$, then $\psi(N_{C(G)}(x))=N_{C(G)}(y)$ since automorphisms preserve adjacency. Since for every $v\in V_{1}$, $\psi(v)=v$, we have $\psi(N_{C(G)}(x))=N_{C(G)}(x)$. Thus, $N_{C(G)}(x)=N_{C(G)}(y)$, which is a contradiction. ∎
@@
In $C(G)$, each vertices of $V_{1}$ is of degree $n-1\geq 3$ and each vertices of $V_{2}$ is of degree $2$. The rest follows from the fact that $deg_{C(G)}(\phi(v))=deg_{C(G)}(v)$ for any $v\in V(C(G))$. ∎
@@
$\pi_{\alpha}(w)=\begin{cases}\alpha(w)&\text{if }w\in V(G)\\ w_{\alpha(x),\alpha(y)}&\text{if }w\in V(C(G))\backslash V(G),w=w_{x,y},x,y\in V(G).\end{cases}$
We can see that $\pi_{\alpha}\in Aut(C(G))$. If $\{x,y\}\in E(C(G))$ then $\{x,y\}\in E(\overline{G})$ or $\{x,y\}\in E(S(G))$. By the definition of $\pi_{\alpha}$, if $\{x,y\}\in E(S(G))$ then $\{\pi_{\alpha}(x),\pi_{\alpha}(y)\}\in E(S(G))$. Moreover, we have $\{x,y\}\in E(\overline{G})\iff\{x,y\}\not\in E(G)\iff\{\alpha(x),\alpha(y)\}\not\in E(G)\iff\{\alpha(x),\alpha(y)\}\in E(\overline{G})\iff\{\pi_{\alpha}(x),\pi_{\alpha}(y)\}\in E(\overline{G})$. Thus, if $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$, then $H\subseteq Aut(C(G))$. We show that $Aut(C(G))\subseteq H$. Let $f\in Aut(C(G))$. By Lemma 2.4, we have $f(V_{1})=V_{1}$ since $G$ is a connected graph of order at least $4$. If $\alpha=f\restriction V_{1}$, then $\alpha\in Aut(G)$. Let $\pi_{\alpha}$ be the induced automorphism of $\alpha$ as defined above. If we let $l=f^{-1}\pi_{\alpha}$, then $l$ is an automorphism of $C(G)$ such that $l(v)=v$ for every $v\in V_{1}$. By Lemma 2.3, $l=id_{C(G)}$. Consequently, $f=\pi_{\alpha}$. Therefore, $Aut(C(G))\subseteq H$, and so $H=Aut(C(G))$. It is trivial to see that $H\cong Aut(G)$, so we are done. ∎
@@
Let $G=(V(G),E(G))$ be a graph where $V(G)=\{v_{1},...,v_{p}\}$ for some integer $p\geq 1$. To $G$, we add $p$ new vertices $\{u_{1},...,u_{p}\}$ and $p$ new edges $\{u_{i},v_{i}\}$. The new graph is the endline graph of $G$ denoted by $G^{+}$. The edges $\{u_{i},v_{i}\}$ are the endline edges of $G^{+}$. Let $U_{1}=V(G)$ and $U_{2}=V(G^{+})\backslash V(G)$. Let $L(G)$ be the line graph of $G$.
Fact 2.7.
@@
    (Mirafzal; [25, Theorem 4]) If $G\not\cong C_{n}$ for any $n\geq 3$, then $Aut(S(G))\cong Aut(G)$.
@@
Let $u_{i}\in U_{2}$. If $\psi(u_{i})=u_{j}$ for some $u_{j}\in U_{2}\backslash\{u_{i}\}$, then $\psi(v_{i})=v_{j}$ since the automorphisms preserve adjacency, which is a contradiction to the given assumption. ∎
@@
In $G^{+}$, each vertex of $U_{1}$ has a degree at least $2$ and each vertex of $U_{2}$ has a degree $1$. ∎
@@
Since $G^{+}\not\in\{Q,L(Q),P_{2}\}$ as $G^{+}$ has $n$ vertices of degree one (see Fig. 1), $Aut(M(G))\cong Aut(L(G^{+}))\cong Aut(G^{+})$ by Fact 2.7. Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(G^{+})\rightarrow V(G^{+})$ by the following rule:
@@
Then $\pi_{\alpha}\in Aut(G^{+})$, where $\pi_{\alpha}(U_{i})=U_{i}$ for $i\in\{1,2\}$. The rest follows by applying Lemmas 2.8, 2.9, and the arguments in the proof of Lemma 2.5. ∎
By Lemmas 2.5, 2.10, and Fact 2.7(2,3), we have the following result.

### v2

# Automorphism groups and Distinguishing Colorings of Central and Middle Graphs

@@ Abstract.
Let $G$ be a simple, finite, connected, and undirected graph. The middle graph $M(G)$ of $G$ is obtained from the subdivision graph $S(G)$ after joining pairs of subdivided vertices that lie on adjacent edges of $G$ and the central graph $C(G)$ of $G$ is obtained from $S(G)$ after joining all non-adjacent vertices of $G$. We show that if the order of $G$ is at least $4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$(as abstract groups) and apply these results to obtain new upper bounds of the distinguishing number and the distinguishing index of $C(G)$ and $M(G)$ inspired by an algorithm by Kalinowski, Pilśniak, and Woźniak from 2016.
@@
Automorphism group, distinguishing coloring, central graph, middle graph, subdivision graph
@@
Corresponding author. In this revised version, all results except Propositions 3.2 and 3.3 were obtained and written by the first author. The ideas underlying Propositions 3.2 and 3.3 originate from an earlier version of the manuscript by all authors, where they appeared in the form of diagrams. The propositions are formalized and proved here by the first author.
@@ 1. Introduction
All graphs considered in this paper are finite, simple, connected, and undirected. Let $G=(V(G),E(G))$ be a connected graph with the vertex set $V(G)$ and the edge set $E(G)$. Let $\Delta(G)$ represent the maximum degree of $G$. Let $C_{n}$ be a cycle graph of $n$ vertices, $P_{n}$ be a path of $n$ vertices, $K_{n}$ be a complete graph of $n$ vertices, and $K_{n,m}$ be a complete bipartite graph with bipartitions of sizes $n$ and $m$. A claw is the graph $K_{1,3}$. The degree of a vertex $v$ in $G$, denoted by $deg_{G}(v)$, is the number of edges that emerge from $v$. We say $G$ is a regular graph if all vertices of $G$ have the same degree. The degree of a regular graph $G$, denoted by $d(G)$, is the same as the maximum degree $\Delta(G)$ and the minimum degree $\delta(G)$. Let $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ for any $v\in V(G)$. The middle graph $M(G)$ and the central graph $C(G)$, introduced by Hamada and Yoshimura [8] and Vernold [18], respectively, are classes of graphs derived from subdivision graphs. These graph constructions have been extensively studied; see, for example, [4, 5, 6, 7, 11, 12, 13, 15].
Definition 1.1.
The subdivision graph $S(G)$ of $G$ is obtained from $G$ by putting a new vertex in the middle of every edge of $G$. For each edge $xy\in E(G)$, introduce a new vertex $w_{xy}$ (called the subdivision vertex corresponding to the edge $xy$) and replace the edge $xy$ by the two edges $xw_{xy}$ and $w_{xy}y$. Thus, $V(S(G))=V(G)\cup\{w_{xy}:xy\in E(G)\},$ and $E(S(G))=\{xw_{xy},\,w_{xy}y:xy\in E(G)\}.$ The middle graph $M(G)$ of $G$ is the graph whose vertex set is $V(G)\cup E(G)$, and two vertices $x,y$ in the vertex set of $M(G)$ are adjacent in $M(G)$ in the case where one of the following holds:
@@
    $x\in V(G),y\in E(G)$, and $x,y$ are incident in $G$.
The central graph $C(G)$ of $G$ with order $n$ and size $m$ is a graph of order $n+m$ and size $n\choose{2}$ + $m$ that is obtained from $S(G)$ by adding the edges in $E(\overline{G})$, where $\overline{G}$ is the complement graph of $G$.
Definition 1.2.
Let $Aut(G)$ be the automorphism group of $G$, $id_{G}$ be the trivial automorphism of $G$, and $f$ be a vertex coloring or an edge coloring or a total coloring (that is, a coloring of its vertices and edges) of $G$. An automorphism $\phi$ of $G$ preserves $f$ if the following holds:
@@
    $f(x)=f(\phi(x))$ for all $x\in V(G)$ if $f$ is a vertex coloring,
@@
    $f(e)=f(\phi(e))$ for all $e\in E(G)$ if $f$ is an edge coloring,
The notions of distinguishing number and distinguishing index were introduced by Albertson and Collins [1] and Kalinowski and Pilsniak [10], respectively.
Definition 1.3.
The distinguishing number (index) of a graph $G$, denoted by $D(G)$ ($D^{\prime}(G)$), is the least integer $d$ such that $G$ admits a vertex coloring (edge coloring) with $d$ colors that is preserved only by $id_{G}$.
1.1. Motivation
Alikhani and Soltani [3, Theorems 3.8 and 4.3] studied the distinguishing number and distinguishing index of subdivision graphs and proved that if $G$ is a connected graph of order $n\geq 3$, then
$D(S(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil\quad\text{and}\quad D^{\prime}(S(G))\leq\min\{r:r^{2}-r\geq D^{\prime}(G)\},$
where $\Delta(G)$ denotes the maximum degree of $G$. Kalinowski–Pilśniak–Woźniak [9] investigated the total distinguishing chromatic number of graphs. Inspired by the results of Alikhani and Soltani [3] and the results of Kalinowski–Pilśniak–Woźniak [9], we investigate the distinguishing parameters of middle and central graphs.
1.2. Main Results
Let $G$ be a connected graph of order $n$. We prove the following:
@@ 1.2. Main Results
    (Theorem 2.9) If $n\geq 4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$.
@@
    (Theorem 3.10(1)) If $n\geq 3$, then $D(M(G))\leq\Delta(G)$;
@@
    (Theorem 3.10(2)) $D^{\prime}(M(G))\leq 3$;
@@
    (Theorem 3.4) If $n\geq 4$, then $D^{\prime}(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil.$
@@
    (Theorem 3.8) If $n\geq 4$, then $D(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil.$
@@ 2. Automorphism groups
In this section, we sketch some lemmas concerning $Aut(C(G))$ and $Aut(M(G))$ where $M(G)$ is the middle graph of $G$. Applying these lemmas, we will prove Theorems 3.4 and 3.8 and propositions 3.2, 3.3, and 3.6. Recall that for each edge $xy\in E(G)$, $w_{xy}$ denotes the subdivision vertex corresponding to the edge $xy$ in the subdivision graph $S(G)$. In $S(G)$, let $V_{1}=V(G)$ be the set of all vertices of $G$, and let $V_{2}=\{w_{xy}:xy\in E(G)\}$ denote the set of all subdivision vertices.
@@
First, we observe that distinct vertices of $V_{2}$ have distinct neighborhoods in $C(G)$. Indeed, if $w_{xy},w_{uv}\in V_{2}$ with $w_{xy}\neq w_{uv}$, then $\{x,y\}\neq\{u,v\}$ since $G$ is simple, and hence each edge of $G$ has a unique pair of endpoints. Therefore,
$N_{C(G)}(w_{xy})=\{x,y\}\neq\{u,v\}=N_{C(G)}(w_{uv}).$
Now, let $x\in V_{2}$. If $\psi(x)=y$ for some $y\in V_{2}\setminus\{x\}$, then $\psi(N_{C(G)}(x))=N_{C(G)}(y)$ since automorphisms preserve adjacency. Since $N_{C(G)}(x)\subseteq V_{1}$ and $\psi(v)=v$ for every $v\in V_{1}$, it follows that $\psi(N_{C(G)}(x))=N_{C(G)}(x)$. Thus, $N_{C(G)}(x)=N_{C(G)}(y)$, a contradiction. Since $\psi(u)=u$ for all $u\in V_{1}$, we have $\psi(V_{1})=V_{1}$, and hence $\psi(x)\notin V_{1}$. Therefore, $\psi(x)=x$ for all $x\in V_{2}$, and consequently $\psi=id_{C(G)}$. ∎
@@
In $C(G)$, each vertex of $V_{1}$ has degree $n-1\geq 3$, whereas each vertex of $V_{2}$ has degree $2$. Since automorphisms preserve vertex degrees, that is, $deg_{C(G)}(\varphi(v))=deg_{C(G)}(v)$ for any $v\in V(C(G))$, it follows that $\varphi(V_{1})=V_{1}$ and $\varphi(V_{2})=V_{2}$. ∎
@@
$\pi_{\alpha}(w)=\begin{cases}\alpha(w)&\text{if }w\in V(G)\\ w_{\alpha(x)\alpha(y)}&\text{if }w\in V(C(G))\backslash V(G),w=w_{xy},x,y\in V(G).\end{cases}$
We can see that $\pi_{\alpha}\in Aut(C(G))$. If $xy\in E(C(G))$ then either $xy\in E(\overline{G})$ or $xy\in E(S(G))$. By the definition of $\pi_{\alpha}$, if $xy\in E(S(G))$ then $\pi_{\alpha}(x)\pi_{\alpha}(y)\in E(S(G))$. Moreover, we have
$xy\in E(\overline{G})$ $\iff xy\notin E(G)$
$\iff\alpha(x)\alpha(y)\notin E(G)$
$\iff\alpha(x)\alpha(y)\in E(\overline{G})$
$\iff\pi_{\alpha}(x)\pi_{\alpha}(y)\in E(\overline{G}).$
Thus, if $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$, then $H\subseteq Aut(C(G))$. We show that $Aut(C(G))\subseteq H$. Let $f\in Aut(C(G))$. By Lemma 2.2, we have $f(V_{1})=V_{1}$ since $G$ is a connected graph of order at least $4$. Thus, if $\alpha=f\restriction V_{1}$, then $\alpha\in Aut(G)$. Let $\pi_{\alpha}$ be the induced automorphism of $\alpha$ as defined above. If we let $l=f^{-1}\circ\pi_{\alpha}$, then $l$ is an automorphism of $C(G)$ such that $l(v)=v$ for every $v\in V_{1}$. By Lemma 2.1, $l=id_{C(G)}$. Consequently, $f=\pi_{\alpha}$. Therefore, $Aut(C(G))\subseteq H$, and so $H=Aut(C(G))$. It is trivial to see that $H\cong Aut(G)$, so we are done. ∎
@@
[…]
Fact 2.5.
@@
    (Mirafzal; [14, Theorem 4]) If $G\not\cong C_{n}$ for any $n\geq 3$, then $Aut(S(G))\cong Aut(G)$.
@@
Let $u_{i}\in U_{2}$. Suppose $\psi(u_{i})=u_{j}$ for some $u_{j}\in U_{2}$ with $j\neq i$. Since automorphisms preserve adjacency and $u_{i}$ is adjacent only to $v_{i}\in U_{1}$, it follows that $\psi(v_{i})=v_{j}$. However, by assumption $\psi(v_{i})=v_{i}$ for all $v_{i}\in U_{1}$, which is a contradiction. Thus $\psi(u_{i})=u_{i}$ for all $u_{i}\in U_{2}$, and hence $\psi=id_{G^{+}}$. ∎
@@
In $G^{+}$, each vertex of $U_{1}$ has degree at least $2$, while each vertex of $U_{2}$ has degree $1$. Since automorphisms preserve vertex degrees, it follows that $\varphi(U_{1})=U_{1}$ and $\varphi(U_{2})=U_{2}$. ∎
@@
Since $G^{+}\not\in\{Q,L(Q),P_{2}\}$ as $G^{+}$ has $n$ vertices of degree one (see Fig. 1), $Aut(M(G))\cong Aut(L(G^{+}))$ by the Fact 2.5(1), and $Aut(L(G^{+}))\cong Aut(G^{+})$ by the Fact 2.5(2). Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(G^{+})\rightarrow V(G^{+})$ by the following rule:
@@
Then $\pi_{\alpha}\in Aut(G^{+})$, where $\pi_{\alpha}(U_{i})=U_{i}$ for $i\in\{1,2\}$. The rest follows by applying Lemmas 2.6, 2.7, and the arguments in the proof of Lemma 2.3. ∎
By Lemmas 2.3 and 2.8, together with Fact 2.5(2,3), we obtain the following result.
@@
Proof.
Since $n\geq 5$, we have $G\not\cong P_{2},Q,$ or $L(Q)$. Hence, by Fact 2.5(2), $Aut(G)\cong Aut(L(G)).$ Since $G\not\cong C_{n}$, Fact 2.5(3) implies that $Aut(S(G))\cong Aut(G).$ Moreover, by Lemma 2.3 and Lemma 2.8, we have $Aut(C(G))\cong Aut(G)\text{ and }Aut(M(G))\cong Aut(G).$